OpenStudy (anonymous):

which of the following equations is written in standard form? 13=-7-3x 3x + 6y =13 20y=20 would it be the 2nd one?

OpenStudy (marissalovescats):

Standard form is: Ax+By=C Where A, B, and C are numbers.

OpenStudy (pinksapphire):

Ax+By=C

OpenStudy (marissalovescats):

So yes, the second one :)

OpenStudy (pinksapphire):

The second equation...

OpenStudy (pinksapphire):

Looks like I copied you lol but you just write fast :)

OpenStudy (anonymous):

it is asking me to rewrite in standard form -3x + 6y=12

OpenStudy (anonymous):

i beleive i got this wrong.

OpenStudy (anonymous):

but thank you guys for helping me :)

OpenStudy (marissalovescats):

-3x + 6y=12 is in standard form

OpenStudy (pinksapphire):

It's already in standard form...

OpenStudy (pinksapphire):

LOL again...

OpenStudy (anonymous):

it wanted to rewrite the equation?

OpenStudy (marissalovescats):

Rewrite it to be in what form though?

OpenStudy (anonymous):

it says rewrite in standard form :O

OpenStudy (marissalovescats):

But it already is in standard form. Do you have answer choices?

OpenStudy (pinksapphire):

-3x+6y=12 +3x +3x 6y=12+3x 6y/6=12+3x/6 y=2+1/2x If you need it in y-intercept form then here you go...

OpenStudy (anonymous):

A. 3x - 6y= 12 B.3X - 6y = -12 C.-x+2y=4 @marissalovescats

OpenStudy (marissalovescats):

-3x + 6y=12 Would be the same as 3x-6y=-2 if you timed everything by -1.

OpenStudy (anonymous):

so it would be C?

OpenStudy (marissalovescats):

No B

OpenStudy (marissalovescats):

Well... I guess C could work too if you divided everything by 3.. idk this question is really weird

OpenStudy (anonymous):

Also this one saysto rewrite in standard form.. but i got 3x-2y=100